# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ Appendix B Proofs of Theoretical Results
For this proposition, all the expectation and variance terms without subscript are conditioning on $\boldsymbol{\eta}$. We keep using the notations $\theta_{j}=\mathbb{E}[\hat{\theta}^{\mathrm{PT}}_{j}]$, $\mu_{j}=\mathbb{E}[\tilde{Z}^{f}_{j}]$, $\tilde{\sigma}^{2}_{j}=\mathrm{Var}[\hat{\theta}^{\mathrm{PT}}_{j}]$ and $\tilde{\gamma}_{j}=\mathrm{Cov}[\hat{\theta}_{j}^{\mathrm{PT}},\tilde{Z}^{f}_{j}]$. For this proposition, additional assumptions are placed on the data generating process (integrated over $\mathbb{P}_{\eta}$). We first show how they translate to moment conditions on the estimators $\hat{\theta}^{\mathrm{PT}}_{j}$ and $\tilde{Z}^{f}_{j}$.
Lemma B.1.
Proof.
By Minkowski’s inequality, we have
$=\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[Y_{ij}^{4}\right]^{1/4}<\infty.$
$\theta_{j}^{4}$ $=\mathbb{E}_{\eta_{j}}\!\left[\hat{\theta}^{\mathrm{PT}}_{j}\right]^{4}\leq\mathbb{E}_{\eta_{j}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{4}\right]$
$\Longrightarrow\quad\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\theta_{j}^{4}\right]$ $\leq\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{4}\right]<\infty,$
With Lemma B.1, we will prove Proposition 5.1 via the following steps.
Step 1: Decompose the difference.
We first decompose both CURE and the loss separately as
$\text{CURE}(\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega})$ $=\frac{1}{m}\sum_{j=1}^{m}\left((2\omega_{j}-1)\tilde{\sigma}^{2}_{j}+\left[(1-\omega_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\tilde{Z}^{f}_{j})\right]^{2}+2(1-\omega_{j})\tilde{\gamma}_{j}\right)$
$=\underbrace{\frac{1}{m}\sum_{j=1}^{m}\left((2\omega_{j}-1)\tilde{\sigma}^{2}_{j}+(1-\omega_{j})^{2}(\hat{\theta}^{\mathrm{PT}}_{j}-\mu_{j})^{2}\right)}_{\mathbb{I}(\omega)}$
$+\underbrace{\frac{1}{m}\sum_{j=1}^{m}\left(2(1-\omega_{j})\tilde{\gamma}_{j}+2(1-\omega_{j})^{2}(\mu_{j}-\tilde{Z}^{f}_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\mu_{j})+(1-\omega_{j})^{2}(\mu_{j}-\tilde{Z}^{f}_{j})^{2}\right)}_{\mathbb{II}(\omega)}$
$\ell_{m}(\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\omega},\boldsymbol{\theta})$ $=\frac{1}{m}\sum_{j=1}^{m}(\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\tilde{Z}^{f}_{j}-\theta_{j})^{2}$
$=\underbrace{\frac{1}{m}\sum_{j=1}^{m}(\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\mu_{j}-\theta_{j})^{2}}_{\mathbb{I}^{*}(\omega)}$
$+\underbrace{\frac{1}{m}\sum_{j=1}^{m}\left(2(1-\omega_{j})(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})+2(1-\omega_{j})^{2}(\mu_{j}-\tilde{Z}^{f}_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\mu_{j})+(1-\omega_{j})^{2}(\mu_{j}-\tilde{Z}^{f}_{j})^{2}\right)}_{\mathbb{II}^{*}(\omega)},$
and we are interested in bounding
The proof in this step is directly adapted from Theorem 5.1 in Xie et al. [2012] and generalizes to non-Gaussian data. With some algebraic manipulation, we can further decompose
$\Delta_{1}(\omega)$ $=\frac{1}{m}\sum_{j=1}^{m}\left((2\omega_{j}-1)\tilde{\sigma}^{2}_{j}+(1-\omega_{j})^{2}(\hat{\theta}^{\mathrm{PT}}_{j}-\mu_{j})^{2}\right)$
$-\frac{1}{m}\sum_{j=1}^{m}(\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\mu_{j}-\theta_{j})^{2}$
$=\text{CURE}(\boldsymbol{\hat{\theta}}^{0}_{\omega})-\ell_{m}(\boldsymbol{\hat{\theta}}^{0}_{\omega},\boldsymbol{\theta})-\frac{2}{m}\sum_{j=1}^{m}\mu_{j}(1-\omega_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})$
$\text{where}\quad\text{CURE}(\boldsymbol{\hat{\theta}}^{0}_{\omega})$ $=\frac{1}{m}\sum_{j=1}^{m}\left((2\omega_{j}-1)\tilde{\sigma}^{2}_{j}+(1-\omega_{j})^{2}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}\right),\quad\ell_{m}(\boldsymbol{\hat{\theta}}^{0}_{\omega},\boldsymbol{\theta})=\frac{1}{m}\sum_{j=1}^{m}(\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})^{2},$
corresponds to CURE and the loss of the “shrink-towards-zero” estimator $\hat{\theta}^{0}_{j,\omega}:=\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}$. We thus have
$\sup_{\omega\geq 0}\left|\Delta_{1}(\omega)\right|$ $\leq\sup_{\omega\geq 0}\left|\text{CURE}(\boldsymbol{\hat{\theta}}^{0}_{\omega})-\ell_{m}(\boldsymbol{\hat{\theta}}^{0}_{\omega},\boldsymbol{\theta})\right|+\frac{2}{m}\sup_{\omega\geq 0}\Big{|}\sum_{j=1}^{m}\mu_{j}(1-\omega_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\Big{|}.$ (22)
Now, rearrangements of terms gives that
$\sup_{\omega\geq 0}\left|\text{CURE}(\boldsymbol{\hat{\theta}}^{0}_{\omega})-\ell_{m}(\boldsymbol{\hat{\theta}}^{0}_{\omega},\boldsymbol{\theta})\right|$ $=\sup_{\omega\geq 0}\bigg{|}\frac{1}{m}\sum_{j=1}^{m}\left(\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\tilde{\sigma}^{2}_{j}-\theta_{j}^{2}-2\omega_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\right)\bigg{|}$
$\leq\underbrace{\bigg{|}\frac{1}{m}\sum_{j=1}^{m}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\tilde{\sigma}^{2}_{j}-\theta_{j}^{2}\bigg{|}}_{(*)}+\underbrace{\sup_{\omega\geq 0}\bigg{|}\frac{1}{m}\sum_{j=1}^{m}2\omega_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}}_{(**)}.$
For the first term ()(*)( ∗ ),
$\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathbb{E}_{\boldsymbol{\eta}}\!\left[\bigg{(}\frac{1}{m}\sum_{j=1}^{m}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\tilde{\sigma}^{2}_{j}-\theta_{j}^{2}\bigg{)}^{2}\right]\right]=\frac{1}{m^{2}}\sum_{j=1}^{m}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}\right]\right]\leq\frac{1}{m}\sup_{j}\mathrm{Var}_{\mathbb{P}_{\eta}}\!\left[{\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}}\right].$
Thus by Jensen’s inequality and iterated expectation:
$\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\bigg{|}\frac{1}{m}\sum_{j=1}^{m}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\tilde{\sigma}^{2}_{j}-\theta_{j}^{2}\bigg{|}\right]\leq\left(\frac{1}{m}\sup_{j}\mathrm{Var}_{\mathbb{P}_{\eta}}\!\left[{\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}}\right]\right)^{1/2}.$ (23)
$\sup_{\omega\geq 0}\bigg{|}\frac{1}{m}\sum_{j=1}^{m}2\omega_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}$ $\leq\max_{1\geq c_{1}\geq...\geq c_{m}\geq 0}\bigg{|}\frac{2}{m}\sum_{j=1}^{m}c_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}.$ (24)
Lemma B.2.
Let $A_{1},\dots,A_{n}$ be real numbers. Then
Proof.
Since $c_{k}\geq c_{k+1}$, each $\alpha_{k}:=c_{k}-c_{k+1}$ is nonnegative, and
$\sum_{k=1}^{n}\alpha_{k}=c_{1}-c_{n+1}\leq 1.$
Hence,
This shows
$c_{1}=c_{2}=\dots=c_{k}=1,\quad c_{k+1}=c_{k+2}=\dots=c_{n}=0.$
Since $1\geq c_{1}\geq\dots\geq c_{n}\geq 0$, we have that
as claimed. ∎
With Lemma B.2 in hand, we have
$\max_{1\geq c_{1}\geq...\geq c_{m}\geq 0}\bigg{|}\frac{2}{m}\sum_{j=1}^{m}c_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}$ $=\max_{1\leq k\leq m}\bigg{|}\frac{2}{m}\sum_{j=1}^{k}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}.$
Let $M_{k}:=\sum_{j=1}^{k}(\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j})$, it is easy to see that $\{M_{k}\}_{k=1}^{m}$ forms a martingale conditional on $\boldsymbol{\eta}$. Therefore, by a standard $L^{2}$ maximal inequality (ref. Theorem 4.4.6 in Durrett [2019]), we have
$\mathbb{E}_{\boldsymbol{\eta}}\!\left[\max_{1\leq k\leq m}M_{k}^{2}\right]\leq 4\mathbb{E}_{\boldsymbol{\eta}}\!\left[M_{m}^{2}\right]=4\sum_{j=1}^{m}\mathrm{Var}_{\eta_{j}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}\right],$ (25)
which then implies
$\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\bigg{(}\sup_{\omega\geq 0}\bigg{|}\frac{1}{m}\sum_{j=1}^{m}2\omega_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}\bigg{)}^{2}\right]$ $\leq\frac{4}{m^{2}}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\max_{1\leq k\leq m}M_{k}^{2}\right]$
@@
$=\frac{16}{m^{2}}\sum_{j=1}^{m}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}\right]\right]$
$\leq\frac{16}{m}\sup_{j}\mathrm{Var}_{\mathbb{P}_{\eta}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}\right]$
$\Longrightarrow\quad\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\sup_{\omega\geq 0}\bigg{|}\frac{1}{m}\sum_{j=1}^{m}2\omega_{j}\Big{(}\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}-\tilde{\sigma}^{2}_{j}\Big{)}\bigg{|}\right]$ $\leq\left(\frac{16}{m}\sup_{j}\mathrm{Var}_{\mathbb{P}_{\eta}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}\right]\right)^{1/2}.$ (26)
Next, we bound the last expression in (22): $\,\frac{2}{m}\sup_{\omega\geq 0}\left|\sum_{j=1}^{m}(1-\omega_{j})\mu_{j}(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\right|$. Note that $(1-\omega_{j})$ is also monotonic in $\tilde{\sigma}^{2}_{j}$, and if we define $M^{\prime}_{k}:=\sum_{j=1}^{k}\mu_{j}(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})$, then $\{M^{\prime}_{k}\}_{k=1}^{m}$ forms another martingale conditioning on $\boldsymbol{\eta}$. Therefore, following the same argument as (24)–(25) gives
$\frac{4}{m^{2}}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\sup_{\omega\geq 0}\Big{|}\sum_{j=1}^{m}(1-\omega_{j})\mu_{j}(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\Big{|}^{2}\right]$ $\leq\frac{4}{m^{2}}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\max_{1\leq k\leq m}{M^{\prime}_{k}}^{2}\right]$
$\leq\frac{16}{m^{2}}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[{M^{\prime}_{m}}^{2}\right]=\frac{16}{m}\sup_{j}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[\hat{\theta}^{\mathrm{PT}}_{j}\right]\mu_{j}^{2}\right]$
$\Longrightarrow\>\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\frac{2}{m}\sup_{\omega\geq 0}\left|\sum_{j=1}^{m}(1-\omega_{j})\mu_{j}(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\right|\right]$ $\leq\left(\frac{16}{m}\sup_{j}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[\hat{\theta}^{\mathrm{PT}}_{j}\right]\mu_{j}^{2}\right]\right)^{1/2}.$ (27)
The upper bounds derived in (23), (26) and (27) establish control on
$\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\sup_{\omega\geq 0}\left|\Delta_{1}(\omega)\right|\right]\leq\frac{4}{\sqrt{m}}\left(\sup_{j}\mathrm{Var}_{\mathbb{P}_{\eta}}\!\left[{\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}}\right]^{1/2}+\sup_{j}\mathrm{Var}_{\mathbb{P}_{\eta}}\!\left[\big{(}\hat{\theta}^{\mathrm{PT}}_{j}\big{)}^{2}-\hat{\theta}^{\mathrm{PT}}_{j}\theta_{j}\right]+\sup_{j}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[\hat{\theta}^{\mathrm{PT}}_{j}\right]\mu_{j}^{2}\right]\right),$
since each term on the right-hand side can be controlled by the fourth-moment conditions established in Lemma B.1, we know that $\Delta_{1}(\omega)$ converges uniformly to zero.
$\Delta_{2}(\omega)$ $=\frac{2}{m}\sum_{j=1}^{m}(1-\omega_{j})\big{[}\tilde{\gamma}_{j}-(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\big{]}.$ (28)
By the same proof logic that has been applied twice above, we now have a function $(1-\omega_{j})$ monotonic in $\tilde{\sigma}_{j}^{2}$, and a martingale $Q_{k}:=\sum_{j=1}^{k}\left[\tilde{\gamma}_{j}-(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\right]$ for $k=1,\ldots,m$ (recall that $\tilde{\gamma}_{j}=\mathrm{Cov}_{\eta_{j}}[\hat{\theta}^{\mathrm{PT}}_{j},\tilde{Z}^{f}_{j}]$). The steps from (24)–(25) follows, and we have

### v2

@@ Appendix C Details on PAS with Unknown Second Moments
C.1 Sample-based Estimators for $\sigma_{j}^{2},\tau_{j}^{2},\gamma_{j}$
We first write down the expressions of the unbiased sample-based estimators for $\sigma_{j}^{2},\tau_{j}^{2}$ and $\gamma_{j}$, assuming that $n_{j}\geq 2$.
$\hat{\sigma}_{j}^{2}$ $:=\frac{1}{n_{j}-1}\sum_{i=1}^{n_{j}}(Y_{ij}-\bar{Y}_{j})^{2},\quad\hat{\tau}_{j}^{2}:=\frac{1}{n_{j}+N_{j}-1}\bigg{(}\sum_{i=1}^{n_{j}}(f(X_{ij})-\bar{Z}^{N+n}_{j})^{2}+\sum_{i=1}^{N_{j}}(f(\tilde{X}_{ij})-\bar{Z}^{N+n}_{j})^{2}\bigg{)},$
These sample-based estimators serve a dual role. Firstly, they are utilized in the practical implementation of the PT and PAS estimators for our numerical experiments on real-world datasets, where true second moments are unavailable. Secondly, they form the basis for defining the UniPT and UniPAS estimators, which we introduce subsequently.
C.2 UniPT: Power-tuning Across Problems with Estimated Second Moments
Under this new setup, the first step is to derive a new variant of the PT estimator without the known second moments. It turns out that performing power-tuning across problems using the same $\lambda$ for all problems (which we show in Section A.1 to be closely related to the multivariate estimation problem in PPI++) leads to a promising alternative.
Definition C.1 (Univariate Power Tuning (UniPT)).
We consider a family of estimators for each problem $j\in[m]$:
$\hat{\theta}_{j,\lambda}=\bar{Y}_{j}+\lambda(\bar{Z}^{f}_{j}-\tilde{Z}^{f}_{j}),$
where all problems are controlled by a single, global power-tuning parameter $\lambda\in\mathbb{R}$. Our goal is to find the $\lambda$ that minimizes the sum of variances across all problems, $\sum_{j=1}^{m}\mathrm{Var}_{\eta_{j}}[\hat{\theta}_{j,\lambda}]$. The variance for each problem $j$ is:
@@ C.2 UniPT: Power-tuning Across Problems with Estimated Second Moments
$\mathrm{Var}_{\eta_{j}}[\hat{\theta}_{j,\lambda}]=\frac{\sigma_{j}^{2}}{n_{j}}+\lambda^{2}\left(\frac{1}{n_{j}}+\frac{1}{N_{j}}\right)\tau_{j}^{2}-\frac{2\lambda}{n_{j}}\gamma_{j},$
where $\sigma_{j}^{2}=\mathrm{Var}_{\eta_{j}}[Y_{ij}]$, $\tau_{j}^{2}=\mathrm{Var}_{\eta_{j}}[f(X_{ij})]$, and $\gamma_{j}=\mathrm{Cov}_{\eta_{j}}[Y_{ij},f(X_{ij})]$. Minimizing $\sum_{j=1}^{m}\mathrm{Var}_{\eta_{j}}[\hat{\theta}_{j,\lambda}]$ with respect to $\lambda$ yields the theoretically optimal global parameter for the given set of $m$ problems:
$\lambda^{*}_{m}:=\frac{\sum_{j=1}^{m}n_{j}^{-1}\gamma_{j}}{\sum_{j=1}^{m}\left(\frac{1}{n_{j}}+\frac{1}{N_{j}}\right)\tau_{j}^{2}}=\frac{\sum_{j=1}^{m}n_{j}^{-1}\gamma_{j}}{\sum_{j=1}^{m}\frac{n_{j}+N_{j}}{n_{j}N_{j}}\tau_{j}^{2}}.$
In practice, since $\gamma_{j}$ and $\tau_{j}^{2}$ are now assumed unknown, we replace them with their sample-based unbiased estimators $\hat{\gamma}_{j}$ and $\hat{\tau}_{j}^{2}$ (computed in Section C.1) to obtain the estimated global power-tuning parameter:
$\hat{\lambda}:=\frac{\sum_{j=1}^{m}n_{j}^{-1}\hat{\gamma}_{j}}{\sum_{j=1}^{m}\frac{n_{j}+N_{j}}{n_{j}N_{j}}\hat{\tau}_{j}^{2}}.$
We clip $\hat{\lambda}$ to the interval $[0,1]$. Let this be $\hat{\lambda}_{\mathrm{\mathrm{clip}}}=\mathrm{\mathrm{clip}}(\hat{\lambda},[0,1])$. Similarly, we denote the clipped version of $\lambda^{*}_{m}$ as $\lambda^{*}_{\mathrm{clip},m}:=\mathrm{\mathrm{clip}}(\lambda^{*}_{m},[0,1])$. The UniPT estimator for problem $j$ is then:
$\hat{\theta}_{j}^{\mathrm{UPT}}=\bar{Y}_{j}+\hat{\lambda}_{\mathrm{clip}}(\bar{Z}^{f}_{j}-\tilde{Z}^{f}_{j}).$
The UniPT estimator, by using a single data-driven $\hat{\lambda}$, offers a practical way to perform power tuning when per-problem moments are unknown. This approach is justified by the following theoretical results, which hold as $m\to\infty$.
Proposition C.2 (Asymptotic Consistency of Clipped Global Tuning Parameter).
Assume that:
  1. 1.
    Sample sizes $n_{j},N_{j}$ are bounded ($2\leq n_{\mathrm{min}}\leq n_{j}\leq n_{\mathrm{max}}<\infty$, $1\leq N_{\mathrm{min}}\leq N_{j}\leq N_{\mathrm{max}}<\infty$).
  2. 2.
    $\mathbb{E}_{\mathbb{P}_{\eta}}[\mathrm{Var}_{\eta_{j}}[\hat{\gamma}_{j}]]<\infty$ and $\mathbb{E}_{\mathbb{P}_{\eta}}[\mathrm{Var}_{\eta_{j}}[\hat{\tau}_{j}^{2}]]<\infty$.
  3. 3.
    $\mathbb{E}_{\mathbb{P}_{\eta}}[\gamma_{j}^{2}]<\infty$, $\>\mathbb{E}_{\mathbb{P}_{\eta}}[(\tau_{j}^{2})^{2}]<\infty,\>\mathbb{E}_{\mathbb{P}_{\eta}}[(\sigma_{j}^{2})^{2}]<\infty$. [^18] Note that this condition is implied by the finite fourth-moment assumption in Proposition 5.1.
  4. 4.
    (Denominator bounded away from 0) there exists some $\varepsilon>0$ such that
    $\lim_{m\to\infty}\mathbb{P}\bigg{[}\bigg{|}\frac{1}{m}\sum_{j=1}^{m}\frac{n_{j}+N_{j}}{n_{j}N_{j}}\tau_{j}^{2}\bigg{|}>\varepsilon\bigg{]}=1$
Then, the clipped estimated global tuning parameter $\hat{\lambda}_{\mathrm{clip}}$ converges in $L^{2}$ to the clipped theoretical optimal global tuning parameter $\lambda^{*}_{\mathrm{clip},m}$:
$\hat{\lambda}_{\mathrm{clip}}-\lambda^{*}_{\mathrm{clip},m}\xrightarrow{L^{2}}0\quad\text{as }m\to\infty.$
The proof is deferred to Section F.5.
This consistency ensures that $\hat{\lambda}_{\mathrm{clip}}$ effectively targets the best single power-tuning parameter for the collection of $m$ problems. Building on this, we can state a result regarding the asymptotic variance of the UniPT estimator.
Theorem C.3 (Asymptotic Variance Optimality of UniPT).
Under the assumptions of Proposition C.2, the sum of variances of the UniPT estimators, $\sum_{j=1}^{m}\mathrm{Var}_{\eta_{j}}[\hat{\theta}_{j}^{\mathrm{UPT}}]$, asymptotically achieves the minimum possible sum of variances within the class of estimators $\mathcal{C}=\{(\hat{\theta}_{j,\lambda})_{j=1}^{m}\mid\hat{\theta}_{j,\lambda}=\bar{Y}_{j}+\lambda(\bar{Z}^{f}_{j}-\bar{Z}^{\prime f}_{j}),\lambda\in[0,1]\}$, in the sense that:
[…]
where the left-hand side is still a random variable with randomness from drawing $\eta_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathbb{P}_{\eta}$.
The proof is deferred to Section F.6.
C.3 UniPAS: CURE and Adaptive Shrinkage with Estimated Second Moments
Once we obtain the UniPT estimator, which is asymptotically optimal within a class of unbiased estimators, our next goal is to imitate the steps in Section 4.2 to apply shrinkage across problems. To do so, we must first revisit the formulation of CURE and see how it depends on the now unknown second-moment parameters. By definition:
[…]
So there are three places where CURE makes use of $\tilde{\sigma}_{j}^{2},\tilde{\gamma}_{j}$, which then depend on second moments of the data generating process: (1) in the definition of PT estimator, (2) in the definition of CURE itself and (3) in determining the localized shrinkage level $\omega_{j}$. We thus make the following modifications to CURE using the sample-based estimators.
  1. 1.
    We first replace the PT estimator (shrinkage source) to the UniPT estimator. Additionally, $\tilde{\sigma}_{j}^{2}$ is now the variance of $\hat{\theta}_{j,\lambda^{*}_{\mathrm{clip},m}}$ (as defined in Definition C.1) and $\tilde{\tau}_{j}$ is its covariance with $\tilde{Z}^{f}_{j}$. We can explicitly write them down as
    $\tilde{\sigma}_{j}^{2}:=\frac{\sigma^{2}_{j}}{n_{j}}+\frac{N_{j}+n_{j}}{N_{j}n_{j}}(\lambda^{*}_{\mathrm{clip},m})^{2}\tau_{j}^{2}-\frac{2}{n_{j}}\lambda^{*}_{\mathrm{clip},m}\gamma_{j}\>,\quad\tilde{\gamma}_{j}:=\lambda^{*}_{\mathrm{clip},m}\frac{\tau_{j}^{2}}{N_{j}}.$
  2. 2.
    For $\tilde{\sigma}_{j}^{2}$ and $\tilde{\gamma}_{j}$ in the definition of CURE, we replace them directly with the sample-based estimators
    $\dot{\sigma}_{j}^{2}:=\frac{\hat{\sigma}^{2}_{j}}{n_{j}}+\frac{N_{j}+n_{j}}{N_{j}n_{j}}{\hat{\lambda}}_{\mathrm{clip}}^{2}\hat{\tau}_{j}^{2}-\frac{2}{n_{j}}\hat{\lambda}_{\mathrm{clip}}\hat{\gamma}_{j}\>,\quad\dot{\gamma}_{j}:=\hat{\lambda}_{\mathrm{clip}}\frac{\hat{\tau}_{j}^{2}}{N_{j}}$
    where $\hat{\sigma}^{2}_{j},\hat{\tau}_{j}^{2},\hat{\gamma}_{j}$ are defined in Section C.1.
  3. 3.
    For $\tilde{\sigma}_{j}^{2}$ in the definition of $\omega_{j}$, we replace it with the following averaging estimators:
    $\check{\sigma}^{2}_{j}$ $:=\frac{\bar{\sigma}^{2}}{n_{j}}+\frac{N_{j}+n_{j}}{N_{j}n_{j}}{\hat{\lambda}}_{\mathrm{clip}}^{2}\bar{\tau}^{2}-\frac{2}{n_{j}}\hat{\lambda}_{\mathrm{clip}}\bar{\gamma},\quad\text{where}$ (25)
    $\bar{\sigma}^{2}$ $:=\frac{1}{m}\sum_{j=1}^{m}\hat{\sigma}_{j}^{2},\;\>\bar{\tau}^{2}:=\frac{1}{m}\sum_{j=1}^{m}\hat{\tau}_{j}^{2},\;\>\bar{\gamma}:=\frac{1}{m}\sum_{j=1}^{m}\hat{\gamma}_{j}$
    are the average sample co-(variances) across all $m$ problems. [^19] The rationale is as follows: In defining $\omega_{j}$ we pretend that $\sigma_{j}^{2},\tau_{j}^{2}$ and $\gamma_{j}$ are the same across all problems and so can be estimated consistently. We emphasize that our theoretical results do not require that these second moments be identical (i.e., it is only a working modeling assumption).
With these modifications, we define a new class of shrinkage estimators based on $\hat{\theta}_{j}^{\mathrm{UPT}}$. For any global shrinkage parameter $\omega\geq 0$, the problem-specific shrinkage weight is now defined as:
$\hat{\omega}_{j}:=\hat{\omega}_{j}(\omega)=\frac{\omega}{\omega+\check{\sigma}^{2}_{j}}.$
The corresponding family of shrinkage estimators for problem $j$ is:
$\hat{\theta}_{j,\omega}^{\mathrm{UPAS}}:=\hat{\omega}_{j}\hat{\theta}_{j}^{\mathrm{UPT}}+(1-\hat{\omega}_{j})\tilde{Z}_{j}^{f}.$
We then define the modified CURE, denoted $\widehat{\mathrm{CURE}}$, by taking into account all the changes above. [^20] We use the hat notation for $\widehat{\mathrm{CURE}}$ to make it explicit that $\widehat{\mathrm{CURE}}$ is not an unbiased estimator of risk in finite samples. However, we will show that it is a consistent estimator of risk asymptotically as $m\to\infty$.